# GQE: Generalized Query Expansion for Enhanced Text-Video Retrieval

### v1

@@ Abstract
In the rapidly expanding domain of web video content, the task of text-video retrieval has become increasingly critical, bridging the semantic gap between textual queries and video data. This paper introduces a novel data-centric approach, Generalized Query Expansion (GQE), to address the inherent information imbalance between text and video, enhancing the effectiveness of text-video retrieval systems. Unlike traditional model-centric methods that focus on designing intricate cross-modal interaction mechanisms, GQE aims to expand the text queries associated with videos both during training and testing phases. By adaptively segmenting videos into short clips and employing zero-shot captioning, GQE enriches the training dataset with comprehensive scene descriptions, effectively bridging the data imbalance gap. Furthermore, during retrieval, GQE utilizes Large Language Models (LLM) to generate a diverse set of queries and a query selection module to filter these queries based on relevance and diversity, thus optimizing retrieval performance while reducing computational overhead. Our contributions include a detailed examination of the information imbalance challenge, a novel approach to query expansion in video-text datasets, and the introduction of a query selection strategy that enhances retrieval accuracy without increasing computational costs. GQE achieves state-of-the-art performance on several benchmarks, including MSR-VTT, MSVD, LSMDC, and VATEX, demonstrating the effectiveness of addressing text-video retrieval from a data-centric perspective.
@@ 1 Introduction
The exponential growth of web videos has intensified the demand for text-video retrieval. As a fundamental task in vision-language learning, text-video retrieval aims to retrieve videos that are semantically similar to a text query, allowing users to quickly find target information. Typically, text-video retrieval models are built by training a similarity function between video and text. This is often achieved by first encoding a text and a video separately into a joint latent space and then apply a distance metric such as the cosine similarity between the text and video embeddings.
It is said that “a picture is worth a thousand words". This information imbalance is even more obvious between video and text. A text generally cannot fully capture the entire contents of a video. Instead, it either describes certain frames/regions of the video, or roughly describes the general content of the video at a coarse level. Therefore, a video corresponds to multiple possible texts.
@@
Figure 1: Motivation. An example video and its text annotations from MSR-VTT [65] dataset. Videos contain much richer information than text. A video can be described by numerous possible text queries, while some of them are missing in the dataset.
When it comes to the text-video retrieval task, the information imbalance manifests in two ways. Firstly, it affects model training. Training a text-video retrieval model on video-text datasets under these imbalanced conditions can be considered as training with incomplete labels. This biases the model toward specific granularities or scenes emphasized by text queries in the training data. In Fig. 1, a sample video from the MSR-VTT dataset [65] shows that descriptions for some scenes of a video might be missing in the dataset. The third scene can be described as “A noble girl is surrounded by two girls floating in the air". The cross-modal matching that should exist there is missing, resulting in incomplete alignment. This issue could potentially lead to retrieval failures with queries like “A noble girl is…". We notice that popular video-text datasets are annotated without particular consideration about the completeness of the text queries, especially scene coverage. For text-video retrieval, it would be favorable to have a video-text dataset with complete text queries.
Secondly, the imbalance issue undermines the retrieval process (testing). With the one-to-many nature, given the same video, successful retrieval can be achieved by varying texts over different granularities or emphasizing different scenes or details of a scene. Consider the text-video retrieval task as a dart game: a text query is like a dart, and a video is like a dartboard. The traditional retrieval process is like throwing a dart (query) to see whether it can correctly hit the target field (video). There are multiple positions in the same field of the dartboard (video) that can be targeted by darts (queries), all scoring equally. Accordingly, can we use multiple queries during retrieval to leverage the imbalanced one-to-many property?
Previous methods address the information imbalance by designing cross-modal interaction mechanisms at different granularities, such as sentence-frame level [23] and word-frame level [58], or designing more dedicated model architectures, such as teacher-student [18]. The core idea is to encourage the incomplete text queries to capture the relevant part of the video. We regard these methods as model-centric approaches, as they primarily concentrate on model architecture design. Although they have produced promising results, information imbalance in the data continues to limit model performance.
In this paper, we propose an approach, called Generalized Query Expansion (GQE), from a data-centric perspective, which is orthogonal to the model-centric approaches. Specifically, GQE addresses the imbalance issue via expanding the text queries in both training and testing.
Firstly, GQE expands the text queries in the training dataset to bridge the data imbalance gap. Intuitively, captioning models can generate additional queries. Previous works [62] used a zero-shot video captioning model to generate captions. Note that our mission is to complete the text queries such that they can cover most scenes of the video. Therefore, we first adaptively segment the video into short clips according to its content. Then we apply zero-shot captioning for each scene, bringing novel information as the captioning model is pre-trained on large-scale image-text data. Compared to vanilla video captioning, this approach balances the distribution of text queries across time spans.
Secondly, GQE expands the text queries during retrieval, including a query generation step and a query selection step. In query generation, we propose leveraging advanced Large Language Models (LLMs) to obtain an extensive set of expanded queries. Although query expansion  [55, 46] (e.g, using synonyms) seems to be a traditional technique, we emphasize that how to utilize LLMs for query expansion is still under-explored. In popular CLIP-based [47] text-video retrieval models, LLMs can be regarded as an enhancement of the original text encoder. In addition, we introduce a novel query selection step based on the expanded queries. Although simply involving these queries brings considerable improvement, it also induces higher computation cost and the risk of modifying the sought content. In the metaphor of a dart game, a premise of successful retrieval is that the multiple darts (queries) should be in different positions within the same field of the dartboard (video). Therefore, we devise a query selection module that selects effective queries among the candidates by considering relevance and diversity. It not only lowers computational costs but also enhances performance, outperforming the use of the entire set of generated queries.
Our main contributions are as follows: (i) We investigated the information imbalance problem in the text-video retrieval task and addressed this problem from a data-centric perspective. (ii) We designed a comprehensive video captioning approach to expand the text queries of videos in the training dataset. (iii) We introduced an effective query expansion strategy for text-video retrieval, enhancing retrieval performance while limiting computational costs. (iv) The proposed approach, called GQE, achieves state-of-the-art results across standard benchmarks, including MSR-VTT [65], MSVD [63], LSMDC [50] and VATEX [61].
@@ 2 Related Work
2.1 Vision-Language Models
Vision-language learning is a fundamental task in multimodal field. In the early stage, efforts primarily focused on specific downstream cross-modal tasks, such as image captioning [67, 56, 3], visual question answering (VQA) [2], and text-image retrieval [28], etc. These tasks typically rely on pre-trained models from single modalities. With the rise of multimodal data, approaches like CLIP [47] and ALIGN [27] have emerged, demonstrating promising results on numerous downstream tasks [38, 40, 32] by pre-training on large-scale image-text pair datasets [52, 9]. More recently, large language models (LLMs) [54, 8] have further boosted the vision-language domain, thanks to their strong language understanding performance. Based on LLMs, multimodal models [69, 4] like BLIP-2 [31], LLaVA [34], can do image captioning and visual question answering within one model. In this paper, we bootstrap our retrieval model from the pre-trained CLIP [47] model, following prior research. BLIP-2 [31] model is adopted as an image captioner for expanding text queries.
2.2 Text-Video Retrieval
In the field of text-video retrieval, early works are mainly built on pre-trained unimodal models [11, 21, 35, 59, 60]. With the extracted features, intra-modal and cross-modal interactions are designed to enable effective feature fusion. With the emergence of pre-trained vision-language models (VLMs), CLIP4Clip [38] and Straight-CLIP [43] use the pre-trained cross-modal model CLIP [47] as a backbone, which is shown to be effective on the text-video retrieval task. However, a gap still exists between image and video data. Subsequent studies have turned to fine-tuning on video-text datasets [68, 5, 39] to adapt CLIP to the video domain and designing more specialized cross-modal attention mechanisms [23, 58, 17, 24]. The advancement of VLMs also inspired researchers to utilize pre-trained models to generate auxiliary captions [62, 68] for text-video retrieval data. Cap4video [62] explores an architecture to make use of video captions in text-video retrieval. CLIP-VIP [68] utilizes frame captions to enable diverse loss designs. Our work differs from them in two ways. First, we use a comprehensive caption generation strategy to capture the video scenes as complete as possible. Second, the query expansion is generalized to both training and testing phases.
2.3 Query Expansion
In this section, we focus on text query expansion, as text is the most commonly used query format. Query expansion is a classical technique aimed at improving retrieval quality in information (document) retrieval, particularly for sparse retrieval systems [20], where queries and documents have limited overlapping terms or features, leading to vocabulary mismatches. The main purpose of query expansion is to minimize the lexical gap between the query and target documents by reducing the vocabulary mismatch. As its core, this technique expands a query into new terms that express the same concept. Early works either relied on lexical knowledge bases [55, 46] or utilized pseudo-relevance feedback [30, 48] from external resources. Recent works use generative models [72, 25, 51] to generate expansion terms through model trained on specifically curated datasets. As large language models (LLMs) [54, 8] advanced, some studies began exploring query expansion by prompting LLMs [57, 26].
In the cross-modal retrieval domain, a few early research efforts attempted to employ lexical knowledge bases [14, 41] or pseudo-relevance feedback [19, 70] for query expansion. In this work, we study how to utilize powerful LLMs for query expansion in the text-video retrieval process. Furthermore, we devise an approach to select effective queries that enhance performance and reduce the computational cost introduced by expansion.
@@ 3 Method
Figure 2: Framework of our approach. The left part shows the process of query expansion of training dataset via a comprehensive video captioning approach, including a video temporal segmentation module and a zero-shot captioner. In the middle, we adopt a dual-encoder model with a text-conditioned video pooling design. The right part illustrates query expansion during testing phase, where a query generation module and a query selection module are devised to enhance the retrieval performance.
@@ 3.1 Model
In text-video retrieval, the goal is to learn a similarity function $s(Q,V)$ for a given text query $Q$ and a video $V$. It aims to maximize the similarity score of positive text-video samples and assign lower similarity for irrelevant pairs. Consistent with previous studies [38, 43, 23, 62], we adopt pre-trained CLIP [47] as the backbone. In the process of adapting pre-trained CLIP from image to the video domain, a commonly employed technique is mean-pooling, which aggregates frame embeddings into a global video embedding. However, since videos are much more expressive than texts, typical text-agnostic aggregation methods that encompass the entire video may encode irrelevant information not accounted for in the input text. Hence, we utilize text-conditioned pooling from the previous work, X-Pool [23].
@@
where the query of the cross-attention module is the text embedding $\mathbf{e}_{t}$, the key and value are both frame embeddings $\mathbf{e}_{v}$. The resulting output $\mathbf{e}_{v|t}$ is an aggregated video embedding conditioned on the text. For the sake of simplicity, we omit certain implementation details, such as layer normalization. For more details, please refer to X-Pool [23].
@@
3.2 GQE: Query Expansion of Training Set
The text-conditioned pooling architecture effectively addresses the information imbalance between the text query and the video by capturing video frames relevant to the text. As a result, irrelevant video information not mentioned in the text will be eliminated. However, imbalance in the dataset still limits the model performance. For instance, information considered irrelevant for text query A might be highly relevant to text query B. If query B is not present in the training dataset, this information is unlikely to be captured during training. In such a scenario, the video embedding will be biased toward the content described by query A in the latent space. During retrieval, query B will struggle to effectively retrieve the video. If the text queries could comprehensively capture all scenes of a video, then the information of the video can be sufficiently and evenly mined. To tackle this problem, we propose expanding the text queries in the training set using a comprehensive video captioning approach with two key modules: a video temporal segmentation module and an image captioning module.
As the core of the video temporal segmentation module, Kernel Temporal Segmentation (KTS) [44] algorithm is adopted. KTS divides the video into non-overlapping temporal segments by analyzing a sequence of frame descriptors and identifying the change points. In our implementation, we use a pre-trained CLIP to extract the frame features. Given a video $V\in\mathbb{R}^{F\times H\times W\times 3}$,
@@ 3.2 GQE: Query Expansion of Training Set
3.3 GQE: Query Expansion of Testing
In traditional text-video retrieval, one text query is used to retrieve the target video. However, under the information imbalance condition, a video can be successfully retrieved by multiple possible text queries. In this section, we introduce query expansion during testing.
Query Generation. During query generation, there is no access to the video content or other potential texts related to the video. We must initiate the process from the single existing text query. Large Language Models (LLMs) are powerful text generation models, especially suited for zero-shot, few-shot, or prompt-based learning. In our work, we design prompts to instruct an LLM to expand the text query. We show in the ablations (see Tab. 6) that LLMs are preferable over manual or rule based methods (e.g. NLTK [6]). Given a text query $Q$, we task the LLM to imagine the visual scene and rewrite the text query by diversifying the sentence structure and word usage  [^1] The detailed prompts will be provided in the Appendix. Formally,
@@ 3.3 GQE: Query Expansion of Testing
Utilizing Generated Queries. Next, we explain how the generated queries aid in the retrieval process. A common approach for query expansion used in document retrieval is to create a new, longer text query by appending additional information to the original query. The advantage is that it does not bring additional computation since it only requires one-time similarity computation. However, in our experiments (see Tab. 6), we discovered that concatenating text queries or mean-pooling the text embeddings had a detrimental effect on performance. Another option is to run the retrieval with each query independently and then aggregate the results (similarity or ranking). Our empirical study shows that aggregating the ranking results using a majority voting strategy yielded the best performance.
Despite the significant performance improvement, there remains an evident drawback: the increased computational cost. For instance, when using 1 original query plus 10 expanded queries, we require 11 times the amount of pairwise similarity computations compared to the previous approach. Therefore, we propose to reduce the number of generated queries by a query selection module.

### v2

# Bridging Information Asymmetry in Text-video Retrieval: A Data-centric Approach

@@ Abstract
As online video content rapidly grows, the task of text-video retrieval (TVR) becomes increasingly important. A key challenge in TVR is the information asymmetry between video and text: videos are inherently richer in information, while their textual descriptions often capture only fragments of this complexity. This paper introduces a novel, data-centric framework to bridge this gap by enriching textual representations to better match the richness of video content. During training, videos are segmented into event-level clips and captioned to ensure comprehensive coverage. During retrieval, a large language model (LLM) generates semantically diverse queries to capture a broader range of possible matches. To enhance retrieval efficiency, we propose a query selection mechanism that identifies the most relevant and diverse queries, reducing computational cost while improving accuracy. Our method achieves state-of-the-art results across multiple benchmarks, demonstrating the power of data-centric approaches in addressing information asymmetry in TVR. This work paves the way for new research focused on leveraging data to improve cross-modal retrieval.
@@ 1 Introduction
The explosion of online video content has generated an unprecedented demand for efficient and accurate text-video retrieval systems. As a key task in vision-language learning, text-video retrieval enables users to find semantically relevant videos based on natural language queries, making it indispensable for video search and recommendation systems. Typically, this task is approached by encoding both text and video into a joint latent space and calculating their similarity using a metric like cosine similarity. Yet, despite advances in the field, text-video retrieval remains fundamentally constrained by an often-overlooked problem: information asymmetry.
Information asymmetry refers to the inherent imbalance between the rich, multi-layered content of videos and their more concise textual descriptions. While videos can capture visual scenes, actions, and interactions in great detail, their associated text queries or captions often capture only fragments of this complexity. For example, in datasets like MSR-VTT (Xu et al., 2016b), some key moments in a video may be under-described or even omitted altogether, as illustrated in Fig. 1. This asymmetry complicates both the training and inference processes, as retrieval models may learn to focus on only specific parts of the video while disregarding other potentially relevant aspects. Moreover, when users issue text queries that fail to capture the full semantics of a target video, retrieval accuracy diminishes. These challenges raise an important question: can we mitigate information asymmetry by enriching textual representations to better match the richness of video content?
Most state-of-the-art approaches to text-video retrieval (Gorti et al., 2022; Wu et al., 2023b; Xue et al., 2022), especially those built on the CLIP architecture (Radford et al., 2021), assume a symmetrical relationship between text and video by projecting both into a shared latent space. However, this assumption fails to account for the vast disparity in information density between these two modalities. Some recent methods, such as T-MASS (Wang et al., 2024a) and UAVTR (Fang et al., 2023), have sought to address this imbalance by enriching text embeddings through stochastic or distribution-based mechanisms, yet they focus on model-centric solutions and largely overlook the underlying data asymmetry. In contrast, we propose a novel, data-centric approach that directly addresses this issue by enhancing the textual representations at both the training and retrieval stages.
@@
Figure 1: Videos contain much richer information than text. A video can be described by numerous possible text queries, while some of them are missing in the data.
In this paper, we introduce a unified text enrichment framework designed to bridge the information asymmetry between text and video. Our method operates on multiple levels to comprehensively enrich textual representations. During training, we generate event-level captions for videos using pre-trained Vision-Language Models (VLMs) and an event-aware segmentation mechanism, ensuring that all semantically significant moments in the video are captured. This process results in a richer, more diverse set of textual representations that more accurately reflect the video’s content. During retrieval, we harness the generative power of Large Language Models (LLMs) to expand and diversify text queries, allowing for better coverage of the semantic complexity of the target video. This two-stage approach ensures that the textual modality can more fully represent the richness of the video, significantly narrowing the information gap.
However, simply generating multiple enriched queries introduces its own challenge: not all queries contribute equally to retrieval success. This realization led us to explore how effectively selecting the most relevant queries can further enhance retrieval performance. A key insight from our work is the introduction of the “Oracle Query" concept, which sheds light on the potential of query selection to significantly boost retrieval performance. By selecting the optimal query from a set of enriched queries — the so-called “Oracle Query" — we observed a dramatic improvement in retrieval accuracy. For example, in the MSR-VTT dataset, the Rank-1 metric increased from 46.7% to 61.4%. While this “oracle" approach requires access to ground truth during query selection (which is impractical in real applications), it reveals a critical opportunity: effectively selecting queries can simultaneously reduce computational cost and enhance retrieval accuracy. To capitalize on this finding, we introduce a novel query selection mechanism, which selects the most relevant and diverse queries without requiring oracle-level information. This mechanism evaluates queries based on their semantic relevance and diversity, ensuring that only the most effective queries are used. By doing so, we achieve a further performance improvement with the reduction in computational cost, further narrowing the gap between text and video representations. Our contributions can be summarized as follows:
- We address the problem of information asymmetry in text-video retrieval from a novel data-centric perspective, enriching text representations to better align with video content.
- We propose a unified text enrichment framework that operates during both the training and retrieval stages, leveraging VLMs for event-level captioning and LLMs for query diversification.
- We introduce the concept of "Oracle Query", demonstrating the performance potential of optimal query selection in narrowing the gap between text and video representations.
- We design a query selection mechanism that optimizes the use of enriched queries, reducing computational costs and improving retrieval accuracy.
- Our method achieves state-of-the-art performance on standard benchmarks such as MSR-VTT (Xu et al., 2016b), MSVD (Wu et al., 2017b), LSMDC (Rohrbach et al., 2015b), and VATEX (Wang et al., 2019), demonstrating its robustness and effectiveness across diverse datasets.
@@ 2 Related Work
Text-Video Retrieval
Early text-video retrieval methods primarily relied on pre-trained unimodal models, leveraging hand-crafted features and modeling cross-modal alignment (Gabeur et al., 2020; Wang et al., 2021a; b; Li et al., 2022) The introduction of vision-language models (VLMs), represented by CLIP Radford et al. (2021), have inspired works such as CLIP4Clip (Luo et al., 2022) and Straight-CLIP (Portillo-Quintero et al., 2021) to build retrieval models by leveraging pre-trained shared latent spaces. Subsequent studies have tried to finetune CLIP on video-text datasets with specialized designs for cross-modal attention mechanisms (Gorti et al., 2022; Wang et al., 2022; Fang et al., 2021; Ibrahimi et al., 2023; Li et al., 2024). VLMs have also inspired works utilizing synthetic data for this task. Cap4video (Wu et al., 2023b) and CLIP-VIP (Xue et al., 2022) focus on novel architectures using global video captions or fine-grained frame captions, while InternVid (Wang et al., 2023b) and HAVTR (Wang et al., 2024b) generate more larger datasets to improve performance. Our approach differs by focusing on data-centric solutions to address information asymmetry problem in text-video retrieval. By comprehensively considering both training and retrieval phase, our approach demonstrates significant improvements without the need for architectural changes or large-scale data augmentation. In addressing information asymmetry, T-MASS (Wang et al., 2024a) model the text embedding in the latent space as a stochastic embedding, enriching it with a flexible semantic range. UAVTR (Fang et al., 2023) proposes a uncertainty-adaptive approach that models each look-up as a distribution matching procedure. However, they do not fully resolve the asymmetry issue in terms of data. Our work takes a novel data-centric perspective to tackle this problem, providing a complementary approach to these models.
Foundation Models
The increasing data and compute give birth to large foundation models (FMs) (Bommasani et al., 2021), including unimodal FMs and multimodal FMs. The most representative unimodal FMs is Large Language Models (LLMs) (Touvron et al., 2023; Brown et al., 2020). By training on a huge amount of text data via next-token prediction, LLMs demonstrate exceptional capabilities in language understanding and generation, instruction following, and emergent ability. The development of LLMs has also boosted the vision-language domain in various ways. On the model side, a series of Multimodal Large Language Models (MLLMs) (Wu et al., 2023a; Bai et al., 2024), such as BLIP-2 (Li et al., 2023a), LLaVA (Liu et al., 2023), have been built based on LLMs, demonstrating remarkable performance on vision-language tasks (Wang et al., 2020; Bai et al., 2021) and video tasks (Tang et al., 2023; Bai et al., 2025; Lin et al., 2024; Fan et al., 2023). On the data side, there are some attempts on “re-writing" (Fan et al., 2024) training text data of CLIP (Radford et al., 2021) using LLMs to upgrade the original CLIP model. In contrast to previous methods that focus solely on text generation, our text enrichment framework utilize both VLMs and LLMs to enrich the textual representations, providing a more diverse and contextually rich set of text-video matches.
Query Expansion
Query expansion has long been used in document retrieval to bridge the gap between sparse queries and rich document content (Formal et al., 2021). Traditional approaches use lexical knowledge bases (Voorhees, 1994) or pseudo-relevance feedback (Lavrenko & Croft, 2017), while modern methods leverage deep generative models (Zheng et al., 2020). Some recent studies explore prompting LLMs for query expansion (Wang et al., 2023a), aiming to diversify the query to mitigate vocabulary mismatch. Our method diverges from traditional query expansion in a key way: we tackle the more challenging cross-modal scenario by using LLMs not just to mitigate lexical gaps, but to enrich the query’s semantic depth and diversity, ensuring the generated queries better align with the rich content of video data. This approach allows us to cover a broader range of possible video descriptions while maintaining high retrieval precision.
Figure 2: Illustration of the unified text enrichment framework. The left part shows the process of enriching text representations of training data via a comprehensive video captioning approach, including an event-aware video temporal segmentation module and a pre-trained captioner. In the middle, we adopt a dual-encoder model with a text-conditioned video pooling design. The right part illustrates text enrichment during retrieval phase, where a query generation module, a query selection module, and an aggregation module work together to enhance the retrieval performance.
@@ 3.1 Model
In text-video retrieval, the goal is to learn a similarity function $s(Q,V)$ for a given text query $Q$ and a video $V$. It aims to maximize the similarity score of positive text-video samples and assign lower similarity for irrelevant pairs. Consistent with previous studies Luo et al. (2022); Portillo-Quintero et al. (2021); Gorti et al. (2022); Wu et al. (2023b), we adopt pre-trained CLIP Radford et al. (2021) as the backbone. In the process of adapting pre-trained CLIP from image to the video domain, a commonly employed technique is mean-pooling, which aggregates frame embeddings into a global video embedding. However, since videos are much more expressive than texts, typical text-agnostic aggregation methods that encompass the entire video may encode irrelevant information not accounted for in the input text. Hence, we utilize text-conditioned pooling from the previous work, X-Pool Gorti et al. (2022).
@@
where the query of the cross-attention module is the text embedding $\mathbf{e}_{t}$, the key and value are both frame embeddings $\mathbf{e}_{v}$. The resulting output $\mathbf{e}_{v|t}$ is an aggregated video embedding conditioned on the text. For the sake of simplicity, we omit certain implementation details, such as layer normalization. For more details, please refer to X-Pool Gorti et al. (2022).
@@
3.2 Unified Text Enrichment
3.2.1 Training Phase
The text-conditioned pooling architecture effectively addresses the information asymmetry between the text query and the video by capturing video frames relevant to the text. As a result, irrelevant video information not mentioned in the text will be eliminated. However, asymmetry in the dataset still limits the model performance. For instance, information considered irrelevant for text query A might be highly relevant to text query B. If query B is not present in the training dataset, this information is unlikely to be captured during training. In such a scenario, the video embedding will be biased toward the content described by query A in the latent space. During retrieval, query B will struggle to effectively retrieve the video. If the text queries could comprehensively capture all scenes of a video, then the information of the video can be sufficiently and evenly mined. To tackle this problem, we propose enriching the text queries in the training set using a comprehensive video captioning approach with two key modules: a video temporal segmentation module and an image captioning module.
As the core of the video temporal segmentation module, Kernel Temporal Segmentation (KTS) Potapov et al. (2014); Lin & Lei (2023) algorithm is adopted. KTS divides the video into non-overlapping temporal segments by analyzing a sequence of frame descriptors and identifying the change points. In our implementation, we use a pre-trained CLIP to extract the frame features. Given a video $V\in\mathbb{R}^{F\times H\times W\times 3}$,
@@ 3.2.1 Training Phase
3.2.2 Retrieval Phase
In traditional text-video retrieval, one text query is used to retrieve the target video. However, under the information asymmetry condition, a video can be successfully retrieved by multiple possible text queries. In this section, we introduce text query enrichment in the retrieval phase.
In this phase, there is no access to the video content or other potential texts related to the video. We must initiate the process from the single existing text query. Large Language Models (LLMs) are powerful text generation models, especially suited for zero-shot, few-shot, or prompt-based learning. Therefore, we design prompts to instruct an LLM to understand the given text description, reason the visual scene, and describe the scene in a diverse way to cover various semantic concepts. The detailed prompts will be provided in the Appendix. Ablations (see Tab. 6) suggest that LLMs are preferable over manual or rule based methods (e.g. NLTK Bird et al. (2009)) thanks to its language modeling and reasoning capability. Formally, given a text query $Q$,
@@ 3.2.2 Retrieval Phase
Next, we explain how the generated queries aid in the retrieval process. One straightforward approach is to create a new, longer text description by appending additional information to the original text. The advantage is that it does not bring additional computation since it only requires one-time similarity computation. However, in our experiments (see Tab. 6), we discovered that concatenating text queries or average-pooling the text embeddings had a detrimental effect on retrieval performance. We hypothesize that the text encoder struggles to understand the long description. Another option is to run the retrieval with each query independently and then aggregate the results (similarity or ranking). By experimenting a series of aggregation strategies, our empirical study shows that aggregating the ranking results using a majority voting strategy yielded the best performance.
3.3 Query Selection Mechanism
Despite the performance improvement, several drawbacks remain. First, not all enriched text queries contribute equally to retrieval success. Some contain noisy information due to LLM hallucination, while others are nearly identical to the original query, bringing no novel information. Additionally, more queries increase computational costs. For example, using 1 original and 10 enriched queries requires 11 times more pairwise similarity computations than the previous approach. These challenges motivate us to explore effective query selection.